# A phase transition for the two-dimensional random field Ising/FK-Ising model

### v1

@@ 6.3 The low-temperature case
In this section, we consider the low-temperature case and fix $p>p_{c}$. The main goal of this subsection is to prove that the good external field in Definition 5.7 has a positive $\mathbb{P}$ probability not depending on its size. The strategy we use in the low-temperature regime is similar to that in the high-temperature regime. The difference is that in the high-temperature case, all the clusters are small such that none of them has a typical fluctuation, so they will contribute together and lead to the fluctuation. However, in the low-temperature case, the maximal cluster itself can lead to a big fluctuation. We start with some definitions as in Definitions 6.6 and 6.7, which was hinted in Definition 5.7. For the sake of convenience, we divide the definition of a good external field into the following three sets.
@@
Recall that $f(x)=\log\cosh(x)$ and $\mathcal{C}_{*}$ is the boundary cluster. We use the notation $\mathcal{H}_{\mathrm{low}}^{1}$ to denote the set of external field such that
@@
For any constant $c>0$, let $\mathcal{H}_{\mathrm{low}}^{2}(c)$ denote the set of external field such that
$\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\geq c$
where $\mathsf{var}_{\phi}$ denotes the variance operator under $\phi_{p,\Omega}^{\mathrm{w},0}$.
@@
${\mathbb{E}}_{X}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\Bigg{\rangle}$ $\leq{\mathbb{E}}_{X}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\Bigg{\rangle}$
@@
We choose $\theta_{0}=1$. Let $\mathcal{B}^{+}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\leq N^{-0.1}$ and $\mathcal{B}^{-}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\geq-N^{-0.1}$. We first control the probability of $\mathcal{B}^{+}$ under $\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}$. For any configuration $\omega\in\mathscr{S}_{\mathrm{low},*}=\{\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}|\mathcal{C}|\leq N^{0.1}\}$, we have
@@
where the inequality comes from $f(x)\leq\frac{x^{2}}{2}$. Combined with the exponential tail bound for $\chi$-square distributions ([27, Lemma 1]), it yields that
@@
where the inequality comes from $f(x)\geq\frac{x^{2}-x^{4}}{2}$. Combined with the exponential tail bound for $\chi$-square distributions([27, Lemma 1]) and Lemma B.5, it yields that
@@
${\mathbb{E}}~{}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}$ $=\big{\langle}{\mathbb{E}}~{}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}-\big{\langle}{\mathbb{E}}~{}\epsilon h_{\mathcal{C}_{*}}\cdot\epsilon h_{\tilde{\mathcal{C}_{*}}}\big{\rangle}$
$=\big{\langle}\epsilon^{2}|\mathcal{C}_{*}|\big{\rangle}-\big{\langle}\epsilon^{2}|\mathcal{C}_{*}\cap\tilde{\mathcal{C}_{*}}|\big{\rangle}\geq C_{1}\epsilon^{2}N^{2}.$ (6.20)
@@
${\mathbb{E}}~{}\Big{[}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{]}^{2}$ $\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}^{2}\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{4}h_{\mathcal{C}_{*}}^{4}\big{\rangle}$
$=12\big{\langle}\epsilon^{4}|\mathcal{C}_{*}|^{2}\big{\rangle}\leq C_{2}\epsilon^{4}N^{4}.$ (6.21)
@@
Combining Lemmas 6.12, 6.14 and  B.1, a similar version of Propositions 6.1 and 6.4 could be proved via the same idea of Section 5.2. Thus, we just state the Proposition here and omit further details:
@@
Fix $T<T_{c}(2)$ and thus that $p>p_{c}$. For any constant $\theta>0$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{w},\epsilon h}-\phi_{p,\Omega}^{\mathrm{w},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$.
@@ 6.4 Criticality of $\alpha(T)$
[11, Lemma 3.4] Let $\Omega$ be an $M$-box such that $\Omega\subset\Lambda_{N}$, $\gamma\in\{-1,1\}^{\partial_{int}\Lambda_{N}}$ be the boundary condition and $g\in\mathbb{R}^{\Lambda_{N}\setminus\Omega}$ be the external field on $\Lambda_{N}\setminus\Omega$. Let $\mathcal{H}^{\gamma,g}_{\emptyset}\subset{\mathbb{R}}^{\Omega}$ denote the collection of the external field $h$ such that

### v2

@@ 5.3 The low-temperature case
In this section, we consider the low-temperature case and fix $p>p_{c}$. The main goal of this subsection is to prove that the good external field in Definition 6.7 has a positive $\mathbb{P}$ probability not depending on its size. The strategy we use in the low-temperature regime is similar to that in the high-temperature regime. The difference is that in the high-temperature case, all the clusters are small such that none of them has a typical fluctuation, so they will contribute together and lead to the fluctuation. However, in the low-temperature case, the maximal cluster itself can lead to a big fluctuation. We start with some definitions as in Definitions 5.6 and 5.7, which was hinted in Definition 6.7. For the sake of convenience, we divide the definition of a good external field into the following three sets.
@@
Recall that $f(x)=\ln\cosh(x)$ and $\mathcal{C}_{*}$ is the boundary cluster. We use the notation $\mathcal{H}_{\mathrm{low}}^{1}$ to denote the set of external field such that
@@
For any constant $c>0$, let $\mathcal{H}_{\mathrm{low}}^{2}(c)$ denote the set of external field such that $\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\geq c$ where $\mathsf{var}_{\phi}$ denotes the variance operator under $\phi_{p,\Omega}^{\mathrm{w},0}$.
@@
${\mathbb{E}}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\Bigg{\rangle}$ $\leq{\mathbb{E}}~{}\Bigg{\langle}\exp\left(\frac{\epsilon|h_{\mathcal{C}_{*}}|}{T}\right)\cdot\prod_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}\cosh\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\Bigg{\rangle}$
@@
We choose $\theta_{0}=1$. Let $\mathcal{B}^{+}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\leq N^{-0.1}$ and $\mathcal{B}^{-}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}|}{2T^{2}}\geq-N^{-0.1}$. We first control the probability of $\mathcal{B}^{+}$ under $\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{w},0}$. For any configuration $\omega\in\mathscr{S}_{\mathrm{low},*}:=\{\max_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}|\mathcal{C}|\leq N^{0.1}\}$, we have
@@
where the inequality comes from $f(x)\leq\frac{x^{2}}{2}$. Combined with Lemma B.5, it yields that
@@
where the inequality comes from $f(x)\geq\frac{x^{2}-x^{4}}{2}$. Combined with Lemma B.5, it yields that
@@
${\mathbb{E}}~{}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}$ $=\big{\langle}{\mathbb{E}}~{}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}-\big{\langle}{\mathbb{E}}~{}\epsilon h_{\mathcal{C}_{*}}\cdot\epsilon h_{\tilde{\mathcal{C}_{*}}}\big{\rangle}=\big{\langle}\epsilon^{2}|\mathcal{C}_{*}|\big{\rangle}-\big{\langle}\epsilon^{2}|\mathcal{C}_{*}\cap\tilde{\mathcal{C}_{*}}|\big{\rangle}\geq C_{1}\epsilon^{2}N^{2}.$ (5.20)
@@
${\mathbb{E}}~{}\Big{[}\mathsf{var}_{\phi}\big{(}\epsilon h_{\mathcal{C}_{*}}\big{)}\Big{]}^{2}$ $\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{2}h_{\mathcal{C}_{*}}^{2}\big{\rangle}^{2}\leq 4{\mathbb{E}}~{}\big{\langle}\epsilon^{4}h_{\mathcal{C}_{*}}^{4}\big{\rangle}=12\big{\langle}\epsilon^{4}|\mathcal{C}_{*}|^{2}\big{\rangle}\leq C_{2}\epsilon^{4}N^{4}.$ (5.21)
@@
Combining Lemmas 5.12, 5.14 and B.1, a similar version of Propositions 5.1 and 5.4 could be proved, and the proof is the same as that in Section 6.2. Thus, we just state the Proposition here and omit further details:
@@
Fix $T<T_{c}(2)$ and thus that $p>p_{c}$. For any $\theta>0$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{w},\epsilon h}-\phi_{p,\Omega}^{\mathrm{w},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-1}\leq\epsilon\leq 2\theta N^{-1}$.
@@ 5.4 Criticality of $\alpha(T)$
[15, Lemma 3.4] Let $\Omega$ be an $M$-box such that $\Omega\subset\Lambda_{N}$, $\gamma\in\{-1,1\}^{\partial_{int}\Lambda_{N}}$ be the boundary condition and $g\in\mathbb{R}^{\Lambda_{N}\setminus\Omega}$ be the external field on $\Lambda_{N}\setminus\Omega$. Let $\mathcal{H}^{\gamma,g}_{\emptyset}\subset{\mathbb{R}}^{\Omega}$ denote the collection of the external field $h$ such that
@@
6 Singularity in strong-disorder regime
This section is devoted to the proof of singularity results in Theorem 1.3. It combines a coarse-graining argument with some near-critical results that have be shown in Section 5. The proof idea is quite similar to Section 3.2, however, careful analysis of the random cluster configuration is involved since the RFFKIM lacks the nice domain Markov property (DMP) possessed by the RFIM.
6.1 Critical-temperature and high-temperature cases
Critical-temperature and high-temperature regimes can be analyzed jointly. The core intuition is that in both cases, random clusters remain bounded in size – equivalently, dual paths are abundant.
Definition 6.1.
  1. (i)
  2. (ii)
  3. (iii)
For any $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n})$, let $\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})=\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\omega)$ be the marginal distribution on the outmost close region. We emphasize that the event $\mathcal{O}=\mathrm{Out}(\omega)$ is measurable under $\omega_{|E(\Lambda_{N})\setminus E(\Omega(\mathcal{O}))}$. We further define $\Phi_{p,\mathcal{O}}^{h}$ to be the product measure of $\phi^{\mathrm{f},h}_{p,\Omega_{i}}~{}(i=1,2,\cdots,n)$. More precisely, let $\omega^{\mathcal{O}}\in\{0,1\}^{E(\Omega(\mathcal{O}))}$ be an edge configuration and $\omega_{i}^{\mathcal{O}}=\omega^{\mathcal{O}}_{|\Omega_{i}},$ then we define $\Phi_{p,\mathcal{O}}^{h}(\omega^{\mathcal{O}})=\prod_{i=1}^{n}\phi_{p,\Omega_{i}}^{\mathrm{f},h}(\omega_{i}^{\mathcal{O}}).$ By applying DMP, we obtain that for any configuration $\omega$ and $\mathcal{O}=\mathrm{Out}(\omega)$,
Similar to (3.5) and (3.6), we apply (6.1) to obtain that
[…]
This naturally leads to the following definition:
Definition 6.2.
$\phi_{p,\Lambda_{N}}^{\gamma,0}\Big{(}\big{\{}\mathcal{O}:\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\geq 1-c\exp(-c^{-1}n)\big{\}}\Big{)}\geq 1-c\exp(-c^{-1}n).$ (6.4)
Lemma 6.3.
Proof.
Recall that $\{B_{i}\}_{1\leq i\leq n}$ is a partition of $\Lambda_{N}$ by $2M$-boxes. Applying the RSW theory (see [22] and also e.g. [21, 36, 29] for other remarkable progress on the RSW theory), there exists a constant $C_{1}>0$ such that $\phi^{\xi,0}_{p,B_{i}}(\Omega_{i}\neq\emptyset)\geq C_{1}$ where $\Omega_{i}$ is the region satisfying Definition 6.1 and $\xi$ is an arbitrary boundary condition on $\partial_{int}B_{i}$. Thus, we obtain by DMP and Lemma B.6 that
Let $\mathcal{A}$ denote the event that $1-\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\leq C_{3}\exp({-}C_{3}^{-1}n)$ with some $C_{3}>0$ to be determined later. Note that $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$ are product measures of $\phi^{\mathrm{f},0}_{p,\Omega_{i}}$ and $\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}$ and Propositions 5.1 and 5.4 show that ${\mathbb{E}}\|\phi^{\mathrm{f},0}_{p,\Omega_{i}}-\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}\|_{\mathrm{TV}}\geq C_{4}>0$. Thus, applying Corollary 3.4 gives that for any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq\frac{C_{1}n}{2}$
Note that we also fix $C_{3}$ here to be the constant given by Corollary 3.4.
$\leq 3c_{1}\exp(-c_{1}^{-1}n).$
The desired result comes from Lemma 6.3. ∎
6.2 The low-temperature case
Now we consider the low-temperature case. Let $M<N$ be a fixed integer with order $O(\epsilon^{-1})$ to be defined later. The main difference from Section 6.1 is that (6.5) does not hold since with high probability there does not exist a dual circuit in $\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i})$. Thus, we need to consider an open circuit instead of a dual circuit, which provides a wired boundary condition instead of a free boundary condition. Further, we emphasize that the second equality in (6.1) fails since the external field outside $\Omega_{i}$ also influences the measure. Nevertheless, the influence of outside external field is manageable if the open circuit is connected to a unique large cluster.
We start with the construction of the desired unique large cluster.
Definition 6.4.
An edge configuration $\omega$ is called well-connected if there exists a unique connected cluster in $\omega_{|\Lambda}$ such that its diameter is not less than $N/2.$ And we will call this cluster the main cluster in $\omega$ for convenience. Besides, we write $\mathcal{W}\subset\{0,1\}^{\mathbb{E}(\Lambda_{N})}$ for the set of well-connected configurations.
Remark 6.5.
We highlight that in the definition of a well-connected configuration $\omega$, the uniqueness of the main cluster is confined to $\omega_{\mid\Lambda}$. In other words, despite this restriction, there remains the possibility that certain small clusters within $\omega_{\mid\Lambda}$ may be connected to a large cluster in $\omega$ that differs from the main cluster. This clarification is solely intended to enhance readers’ understanding and does not impact the proof in any way.
Definition 6.6.
  1. (i)
    For all $i=1,2,\cdots,n,$ either $\Omega_{i}=\emptyset$ or $\Lambda_{M/2}(u_{i})\subset\Omega_{i}\subset\Lambda_{M}(u_{i})$ such that the circuit boundary of $\Omega_{i}$ is open and it is connected to the main cluster of $\omega.$
  2. (ii)
    If $\Omega_{i}\neq\emptyset$, then $\epsilon h$ is $c$-good with respect to $\Omega_{i}$ where the $c$-good external field will be defined in the Definition 6.7 below.
  3. (iii)
    There does not exists a subset $\Lambda_{M/2}(u_{i})\subset\Omega^{\prime}_{i}\subset\Lambda_{M}(u_{i})$ satisfying (i) and (ii) such that $\Omega_{i}\subsetneqq\Omega^{\prime}_{i}$.
  4. (iv)
    Let $\Gamma=\cup_{i=1}^{n}(B_{i}\setminus\Omega_{i})$, then $\tilde{\omega}=\omega_{|\Gamma}$.
Figure 2: An illustration of the outmost good open region. The dotted squares are $\Lambda_{M}(u_{i})$ where $\Lambda_{2M}(u_{i})=B_{i}$. We do not draw the box $\Lambda_{M/2}(u_{i})$, but each red and orange region contains a box of such size. The red color denotes that $h$ is $c$-good with respect to the region, and the orange color denotes that $h$ is not $c$-good with respect to the region.
We use the notation $\mathcal{O}=\mathrm{Out}(\omega,h)$ to denote that $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})$ is the outmost good open configuration of $\omega$. Note that for $\mathcal{O}=\mathrm{Out}(\omega,h)$, $\omega$ is well-connected if and only if $\tilde{\omega}$ is well-connected. Thus, we get that $\{\omega:\mathrm{Out}(\omega,h)=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})\}=\tilde{\omega}\otimes\{0,1\}^{E(\cup_{i=1}^{n}\Omega_{i})}$. Let $\eta(\mathcal{O})$ denote the number of nonempty regions $\Omega_{i}$. We define $\Phi_{p,\mathcal{O}}^{\epsilon h}$ to be FK measure $\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\cdot\mid\tilde{\omega})$ conditioned on $\omega_{|\Gamma}=\tilde{\omega}$. By calculations in (6.3), it suffices to prove a lower bound for $\|\Phi^{\epsilon h}_{p,\mathcal{O}}-\Phi^{0}_{p,\mathcal{O}}\|_{\mathrm{TV}}$ as in (6.2). Unlike Section 6.1, $\Phi^{\epsilon h}_{p,\mathcal{O}}$ is not a product measure of measures supported on $\Omega_{i}$ because the external field outside $\Omega_{i}$ will influence the measure on $\Omega_{i}$. Therefore, we introduce the following definition of good external field.
Definition 6.7.
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
Lemma 6.8.
There exists a constant $\theta_{0}>0$ such that for any $0<\theta<\theta_{0}$ and $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$, we have the following. For any region $\Lambda_{M}\subset\Omega\subset\Lambda_{2M}$ and any $c$-good external field $\epsilon h$ with respect to $\Omega$, let $X^{\pm}=\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\pm\frac{\epsilon h_{\mathcal{C}_{*}}}{T}$. Then we have
$\langle X^{\pm}\rangle^{\mathrm{w},0}_{p,\Omega}\leq\ln\big{(}\langle\exp(X^{\pm})\rangle^{\mathrm{w},0}_{p,\Omega}\big{)}-c_{1}$
for some $c_{1}>0$ depending only on $\theta$ and $c$.
Definition 6.9.
Lemma 6.10.
Lemma 6.11.
For any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq c_{1}n$, we have
where $c_{2}>0$ does not depend on $\mathcal{O}$ or $\epsilon h$.
Proof.
Fix any external field $\epsilon h$ and $\mathcal{O},$ let $I$ denote the set of indices such that $\Omega_{i}\neq\emptyset$ (thus $\epsilon h$ is good with respect to $\Omega_{i}$). For any $\mathrm{Out}(\omega,h)=\mathcal{O}$ and $i\in I$, let $\mathfrak{C}^{i}$ be the collection of clusters in $\omega_{\mid\Omega_{i}}$ and let $\mathcal{C}_{*}^{i}$ be the boundary cluster in $\Omega_{i}$. Let
Recall Definitions 6.4 and 6.6, we get that $\mathcal{C}_{*}^{i}$ are connected by the main cluster of $\omega$. Then we can compute the Radon-Nikodym derivative between $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$
$\langle X_{i}^{\pm}\rangle^{0}_{p,\Omega_{i}}\leq\ln\left(\langle\exp(X_{i}^{\pm})\rangle^{0}_{p,\Omega_{i}}\right)-C_{1}$
Combining (6.11) with (6.9) shows that for any $\omega\in(\mathscr{S}^{+})^{c}\cap(\mathscr{S}^{-})^{c}$
The desired result comes from combining (6.12) with (6.10). ∎
Finally, we come back to the proof of Lemmas 6.8 and 6.10.
Proof of Lemma 6.8.
$\hskip 50.00008pt-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}(\omega_{2})\setminus\{\mathcal{C}_{*}(\omega_{2})\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}(\omega_{2})|}{2T^{2}}\Bigg{|}-\Bigg{|}\frac{\epsilon^{2}(\big{|}|\mathcal{C}_{*}(\omega_{1})|-|\mathcal{C}_{*}(\omega_{2})|\big{|})}{2T^{2}}\Bigg{|}\geq\frac{C_{2}}{2}.$
$\langle X^{+}\rangle^{0}_{p,\Omega}\leq\ln\big{(}\langle\exp(X^{+})\rangle^{0}_{p,\Omega}\big{)}-C_{5}.$
By symmetry, the same result holds for $X^{-}$. ∎
Proof of Lemma 6.10.
For a box $B_{i}=\Lambda_{2M}(u_{i}),$ let $\mathcal{G}(u)$ denote the event that there exist two open circuits in the annuli $\Lambda_{3M}(u)\setminus\Lambda_{2M}(u)$ and $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$ respectively, and these two circuits are connected in $\omega{{}_{\mid\Lambda_{3M}(u)\setminus\Lambda_{M}(u)}}.$ Then by the RSW theory (see [22] for example), there exists a constant $C_{1}>0$ relying only on $p$ such that for any boundary condition $\gamma$ on $\partial_{int}\Lambda_{3M}(u_{i}),$ $\phi^{\gamma,0}_{p,\Lambda_{3M}(u_{i})}(\mathcal{G}(u))\geq 1-\exp(-C_{1}M).$ Therefore, we define $\mathcal{Q}(\omega)=\left(Q_{x}=\mathbf{1}_{\mathcal{G}(2Mx)}:x\in\Lambda_{\frac{N}{2M}}\right)$ to be a random vertex configuration on the box $\Lambda_{\frac{N}{2M}},$ and then by [32], we get that $\mathcal{Q}$ dominates a Bernoulli site percolation $\mathbb{P}_{q},$ where $q$ can be arbitrary close to 1 as $M$ goes to infinity. We define $\mathcal{A}$ to be the event that the following holds:
(1) There exists a unique cluster $C_{\mathcal{Q}}$ in $\mathcal{Q}$ with diameter at least $\frac{N}{4M}$ and volume at least $\frac{C_{2}N^{2}}{M^{2}}$ in $\mathcal{Q}(\omega)$.
(2) For any connected subset $S$ in $\Lambda_{\frac{N}{2M}}$ with diameter at least $\frac{N}{100M}$, we have $S\cap C_{\mathcal{Q}}\neq\emptyset.$
Next we will prove for any $\omega\in\mathcal{A}$, we have $\omega$ is well-connected. It is easy to see that $C_{\mathcal{Q}}$ induces a cluster $\mathcal{C}$ with diameter at least $N/2$ in $\Lambda=\bigcup_{i=1}^{n}(\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i}))$. So it suffices to prove that this is the unique cluster in $\Lambda$ with diameter at least $N/2$. For any cluster $\mathcal{C}^{\prime}$ in $\Lambda$ with diameter at least $N/2$, let $S=\{x\in\Lambda_{\frac{N}{2M}}:\Lambda_{M}(2Mx){\cap\mathcal{C}^{\prime}\neq\emptyset}\}$. By (2), we get that there exists $x\in S\cap C_{\mathcal{Q}}$. Since $Q_{x}=1$ and $x\in{S}$, $\mathcal{C}^{\prime}$ must intersect with the circuit in $\Lambda_{2M}(u)\setminus\Lambda_{M}(u)$. Thus we get $\mathcal{C}^{\prime}$ intersects with $\mathcal{C}$ and $\omega$ is well-connected. By standard percolation result, under the measure $\mathbb{P}_{p},$ $\mathcal{A}$ has probability $1-\exp(-C_{3}\frac{N}{M}).$ Thus, we get $\phi^{\gamma,0}_{p,\Lambda_{N}}(\omega\text{ is well-connected})\geq 1-\exp\left(-C_{4}\frac{N}{M}\right)$ by noticing that a connected component in $\mathcal{Q}$ induces an open cluster in $\omega.$
Figure 3: Illustration of the event $\mathcal{A}$. The blue circuit lies in $\Lambda_{2M}(x)\setminus\Lambda_{M}(x)$, the red circuit in $\Lambda_{3M}(x)\setminus\Lambda_{2M}(x)$, and the green path(s) connect these two circuits. Neighbouring red circuits intersect with each other.
We choose $\theta=1$ and $M$ such that $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$. Furthermore, we choose $c$ in Definition 6.7 to be the first constant in Lemma 5.14. Then by the results in Section 5.3 (see Lemmas 5.12, 5.13 and 5.14), we get that the external field on the region $\Omega$ is good with a positive probability. Hence by Lemma B.6, we get that
and the desired result comes from Markov’s inequality. ∎